OpenStudy (anonymous):

What is the sum of a 6–term geometric series if the first term is 7 and the last term is 229,376?

OpenStudy (anonymous):

good lord

OpenStudy (anonymous):

Yeah :/

OpenStudy (anonymous):

ok we can do it

OpenStudy (anonymous):

7 + 7r+7r^2+7r^3+7r^4+7r^5\] is what you are looking for, and you know that \[7r^5=229,376\] so \[r^5=\frac{229376}{7}=32768\] and so \[r=\sqrt[5]{32768}=8\]

OpenStudy (anonymous):

so we can add them directly or by the formula \[a+ar+ar^2+ar^3+ar^4+ar^5=a(1+r+r^2+r^3+r^4+r^5)=a\frac{r^6-1}{r-1}\]

OpenStudy (anonymous):

\[a\frac{r^6-1}{r-1}\]

OpenStudy (anonymous):

should get \[\frac{7(8^6-1)}{8-1}=8^6-1=262143\] maybe i made a mistake.

OpenStudy (anonymous):

No I think thats right wow thank you so much satellite :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

YOU are the best :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

I'm serious lol :)